The following are numbers obtained from the population samples: 2, 4, 7, 12, 15. Find the standard deviation of the sample.

1
Expert's answer
2022-01-13T12:04:39-0500

n=5

Mean

"\bar{x} = \frac{2+4+7+12+15}{5}= 8"

Sample standard deviation

"s = \sqrt{\frac{(2-8)^2+(4-8)^2+(7-8)^2 +(12-8)^2 +(15-8)^2}{5-1}} \\

= \sqrt{\frac{36 + 16 + 1 + 16 + 49}{4}} \\

= \sqrt{29.5} \\

= 5.43"